OpenStudy (anonymous):

write the result in standard form a+bi. (-4+2i)(7+i)

OpenStudy (anonymous):

can you mutlitply two binomials?

OpenStudy (anonymous):

\[(a+b)(c+d)=ac+ad+bc+bd\]

OpenStudy (anonymous):

2i^2+10i-28

OpenStudy (anonymous):

ok and what is i^2?

OpenStudy (anonymous):

Remember: \[i^2 = -1\]

OpenStudy (anonymous):

-16+5i

OpenStudy (anonymous):

you got it!